OpenStudy (unicwaan):

Can someone help me verify the identity? sine of x divided by one minus cosine of x + sine of x divided by one minus cosine of x = 2 csc x

OpenStudy (btaylor):

so, this? \[\frac{\sin x}{1-\cos{x}} + \frac{\sin x}{1-\cos{x}} = 2\csc x\]

OpenStudy (unicwaan):

yes except the it's 1- cos and 1 + cos sorry

OpenStudy (welshfella):

first convert the 2 fractions on left side to one fraction

OpenStudy (btaylor):

You want to use the identity \(1-\cos^2 x = \sin^2 x\). Multiply the 1 - cos x by (1 + cos x) on top and bottom, and you will have sin^2 x on the bottom.

OpenStudy (unicwaan):

Okay I understand that step BTaylpr but I don't know how the make the other fraction the same like what welshfella wants me to do

OpenStudy (misty1212):

HI!!

OpenStudy (btaylor):

\[\frac{\sin x}{1-\cos{x}} \cdot \frac{1+\cos x}{1+\cos x} + \frac{\sin x}{1+\cos{x}} \cdot \frac{1-\cos x}{1-\cos x} = 2\csc x\]

OpenStudy (misty1212):

you can also just add, you will get it that way too

OpenStudy (unicwaan):

Oh the conjugate! I forgot about that!

OpenStudy (misty1212):

\[\frac{b}{1+a}+\frac{b}{1-a}=\frac{b(1-a)+b(1+a)}{(1-a)(1+a)}\]

OpenStudy (misty1212):

you get \[\frac{2b}{1-a^2}\]

OpenStudy (unicwaan):

Thank you I justed needed that step to get me started, I can figure the rest out :)

OpenStudy (unicwaan):

just*

OpenStudy (btaylor):

no problem